Please walk me through this problem step-by-step:

The average time to serve a customer at a pizza store has a specification of 5 +/- 0.25 minutes. The loss resulting from an unhappy customer is $30.

A. What is the value of (K) in the Taguchi loss function?

B. If the process has an average customer service time of 6 minutes and the standard deviation of the service time is 1 minute, what is the expected loss per customer using the Taguchi loss function?
Answer #1

Average time - 52 0.25 min. Loss = $30 Taguchi loss function t= k14-12 where 55+0.25 30 = k (5.25-5)2 30 = K (0.25) k= 30 .06